OpenStudy (anonymous):

Which of the following is equivalent to 2^4 = 16? a. log2^4 = 16 b. log4^16 = 2 c. log2^16 = 4 d. log4^2 = 16

OpenStudy (lgbasallote):

note that \[\LARGE \log_a b = x \implies a^x = b\]

OpenStudy (anonymous):

Okay, can you talk me step by step on how to work it out?

OpenStudy (lgbasallote):

2^4 = 16 is in the form a^x = b agree?

OpenStudy (anonymous):

Yeah

OpenStudy (lgbasallote):

can you tell me which is a?

OpenStudy (anonymous):

a is 2.

OpenStudy (lgbasallote):

which is x?

OpenStudy (anonymous):

x is 4

OpenStudy (lgbasallote):

and b?

OpenStudy (anonymous):

16

OpenStudy (lgbasallote):

so a = 2 x = 4 b = 16 now can you write it as \[\LARGE \log_a b = x?\]

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

so the answer would be A.

OpenStudy (lgbasallote):

no check again...be sure to sub right

OpenStudy (anonymous):

i thought i did the math right? can you work the problem out for me, so i can see where i went wrong.

OpenStudy (lgbasallote):

look again...a = 2 b = 16 x = 4 \[\Large \log_a b = x\] no math needed

OpenStudy (anonymous):

SO ITS C.

OpenStudy (lgbasallote):

YES!!! YAY

OpenStudy (anonymous):

thanks :)